# CoLoR-Filter: Conditional Loss Reduction Filtering for Targeted Language Model Pre-training

### v1

@@ 1 Introduction
The content of the data that a language model is trained on can have profound effects on its performance and the efficiency of the training process (Rae et al., 2021; Penedo et al., 2023; Soboleva et al., 2023). But it remains an open research question how to decide which data to include in the training set. In this paper, we analyze a family of loss-based approaches for targeted selection of pre-training data, propose a simple approach that outperforms existing methods, and provide some preliminary evidence of favorable scaling properties.
@@
Taken together, our results demonstrate the feasibility of scalable data selection for language modeling. We show that lightly curated web-scraped datasets like C4 contain small subsets that can yield better performance on particular downstream tasks and that with CoLoR-Filter we can find these subsets efficiently.
@@ 5.2 Results
Figure 2: Scaling of final performance with $\tau$ when targeting Books with 150m parameter models. CoLoR-Filter scales best with $\tau$.
@@ 6.2 Results
Note, we also conduct a few more experiments and ablations in the appendix: Appendix E considers using CoLoR-Filter in-distribution to target C4 loss, Appendix F considers applying CoLoR-Filter batchwise rather than globally, Appendix G considers finetuning on $D_{\text{down}}$ after targeted pre-training, and Appendix I inspects some of the selected and excluded examples.
@@ Appendix G Finetuning after targeted pre-training
Table 7: Held out performance after finetuning on downstream data for different pre-trained 150m models. Note that the Random (3.1b tokens) model is equivalent to the conditional model used to select data with CoLoR-Filter ($\tau=64$).

### v3

@@ 1 Introduction
The content of the data that a language model is trained on can have profound effects on its performance and the efficiency of the training process (Rae et al., 2021; Longpre et al., 2023; Penedo et al., 2023; Soboleva et al., 2023; Li et al., 2024). But it remains an open research question how to decide which data to include in the training set. In this paper, we analyze a family of loss-based approaches for targeted selection of pre-training data, propose a simple approach that outperforms existing methods, and provide some preliminary evidence of favorable scaling properties.
@@ 5.2 Results
Figure 2: Scaling of final performance with $\tau$ when targeting Books with 150m parameter models.
@@
Figure 3: Scaling CoLoR-Filter with $\tau$ when training 1.2b models with data selected by 150m models. Curves end when we exhaust the data in C4.
@@ 6.2 Results
Task generalization.
We can also test task generalization beyond the 8 tasks that were used to select the data on a few more tasks that test common sense reasoning (Wang et al., 2019; Socher et al., 2013; Talmor et al., 2018; Sap et al., 2019). Results are presented in Table 2 compared to a random model trained on 10x as much data. The performance indicates that the data selected by CoLoR-Filter are not overfit to the particular evaluation tasks, but captures some general notion of good data for a range of tasks.
Table 2: Task generalization for the 1.2b models with $\tau=64$.
Method copa rte cb sst2 commonsense qa social iqa
Random (25b tokens) 69.2 48.9 42.8 46.8 33.7 42.9
CoLoR-Filter ($\tau=64$, 2.5b tokens) 65.8 52.6 46.0 55.8 32.6 42.7
Note, we also conduct a few more experiments and ablations in the appendix: Appendix E considers using CoLoR-Filter in-distribution to target C4 loss, Appendix F considers applying CoLoR-Filter batchwise rather than globally, Appendix G considers finetuning on $D_{\text{down}}$ after targeted pre-training, Appendix I inspects some of the selected and excluded examples, and Appendix J compared to FineWeb-edu (Penedo et al., 2024).
@@ Appendix G Finetuning after targeted pre-training
Table 8: Held out performance after finetuning on downstream data for different pre-trained 1.2b models.
@@ I.2 Representative examples
Appendix J Comparison to Fineweb-Edu
Concurrent to our initial work, Penedo et al. [2024] released FineWeb-edu, a classifier for educational content that can filter the FineWeb dataset. Here we provide a comparison between CoLoR-Filter and this classifier-based approach.
Specifically, we re-implement the CoLoR-Filter pipeline on top of the Fineweb dataset and with slightly smaller auxiliary models (125m) to make a more fair comparison to FineWeb-edu. Then we compare on the same suite of 8 downstream tasks over various settings of $\tau$ using the two scores: CoLoR-Filter or the FineWeb-edu classifier. We then train larger models (680M parameters) for 10B tokens of selected data. Results are shown in fig. 18. We find that CoLoR-Filter consistently outperforms FineWeb-edu, which is not so surprising since we are doing more targeted data selection by specifically targeting the downstream NLP tasks rather than a general notion of “educational content”.
Figure 18: A comparison of the performance of 680m models trained on 10B tokens selected with various $\tau$ between CoLoR-Filter and FineWeb-edu.
Appendix K Broader Impact
The development of the CoLoR-Filter for data selection has notable broader impacts on both machine learning and society. It enhances efficiency in language model training, leading to reduced computational resources and environmental footprint, while its scalability democratizes access to high-performing models. The method’s success in diverse downstream tasks promises advancements in fields like medical text processing and legal analysis. However, it also raises concerns about dataset bias, necessitating continuous evaluation and updates. Future research should focus on ensuring models do not inherit biases from the selected training data, extending applications, improving efficiency, and implementing safeguards to maximize societal benefits while minimizing risks.
Appendix L Compute resources
All training is conducted on an internal cluster using H100 GPUs. On one GPU, each 150m training run for 3.1b tokens takes about 4 hours, running the auxiliary models offline and in parallel can be faster. Training the 1.2b model to completion takes about 2 days on 4 GPUs.